# Dexterous Contact-Rich Manipulation via the Contact Trust Region

### v1

@@ 8       $n\leftarrow n+1$
Algorithm 1 CTR Trajectory Optimization (CtrTrajOpt)
IV-B Initial Guess Heuristic
As the trajectory optimization problem (36) is nonconvex, the solution that Algorithm 1 converges to can be sensitive to the choice of initial guesses for the input trajectory $\bar{u}_{0:T-1}$. The most uninformed initial guess would be to set the position command to be the current configuration of actuated bodies ($\bar{u}_{t}=\bar{q}^{\mathrm{a}}_{0},\forall t$). However, this can be problematic especially when the robots are not in contact with the object in $q_{0}$, as $\mathbf{B}_{\kappa}^{\mathrm{u}}$, the gradient of object motion w.r.t. robot actions, can get close to 00 even with dynamics smoothing when the distance between them are far away.
@@ IV-B Initial Guess Heuristic
and simulate forward using Drake [53] until the robot is in contact with the object.
IV-C Examples
@@
Applying the initial guess heuristics in Section IV-B brings the robot into contact with the object, yielding ${}^{1}\bar{u}=0.0$ (magenta square in Figure 6). The updated ${}^{1}\bar{u}$ produces the new linearized dynamics (magenta line in Figure 6a), whose slope more closely matches the “in contact” piece of the non-smooth contact dynamics. Meanwhile, the dual feasibility constraint (27d) keeps the ball from pulling the box back.
@@
  • Red zone: If $q^{\mathrm{o}}_{\text{goal}}$ is in the red zone, the optimizer would move backwards but stop at the boundary of the red region, which corresponds to the leftmost point in the RA-CTR in Figure 6c. Without dual feasibility (27d), a naive linear model with an ETR will move backwards into the red region to attempt pulling the object.
@@
Lastly, we plot the A-CTR and the RA-CTR for the above sequence of $\bar{u}$ in Figure 7e. Here we use $\mathbf{\Sigma}=\mathrm{diag}(0.05^{-2},0.05^{-2})$, which bounds $\delta u$ in a circle of radius $0.05$. The curved part of the trust region boundaries arises from the ellipsoidal constraint (25c), whereas the straight part follows from the feasibility constraints (25d) and (25e). Compared with the RA-CTR at ${}^{0}\bar{u}$ (dark yellow), the RA-CTR at ${}^{1}\bar{u}$ (magenta), obtained from applying the initial guess heuristics to ${}^{0}\bar{u}$, approximates the sticking region much better. Furthermore, as ${}^{2}\bar{u}$ (dark green) gets deeper into the “sticking” mode, the straight edge of its RA-CTR aligns more closely with the boundary between sticking and sliding. Finally, compared to the RA-CTRs, the non-relaxed A-CTRs are always smaller. As a result, Algorithm 1 may need more iterations to converge if the RA-CTR constraint in (37d) is replaced by A-CTRs.
IV-D Model Predictive Control
The CTR-based trajectory optimizer in Section IV-A can be readily turned to a controller through Model Predictive Control (MPC). As shown in Algorithm 2, we obtain the optimal trajectory $u^{\star}_{t}$ by solving (36) with Algorithm 1 (Algorithm 2), only use the first input $u^{\star}_{0}$ to deploy on the true CQDC dynamics Algorithm 2, then re-plan after observing the next configuration. Moreover, we initialize $\mathbf{CtrTrajOpt}$ from the initial guess heuristics (Section IV-B) at the first iteration (Algorithm 2), and from the solution of the previous iteration at later iterations (Algorithm 2).
@@ 13       $q_{t+1}=f(q_{t},u_{0}^{\star})$;
V Local MPC Experiments
In this section, we investigate evaluate the performance of the CTR-based MPC in Algorithm 2 under the CQDC dynamics constraints. We are particularly interested in answering the following questions:
@@ V Local MPC Experiments
  • the planar IiwaBimanual system in Example 6, which comprises of 3 unactauted DOFs, 6 actuated DOFs and 29 collision geometries. The whole system is constrained to the $xy$ plane, with gravity pointing along the negative $z$ direction. The bucket measures $0.28\mathrm{m}$ in diameter. The task is to rotate the object, a cylindrical bucket, to target $SE(2)$ poses.
  • the 3D AllegroHand system, which comprises 6 unactauted DOFs, 16 actuated DOFs and 39 collision geometries. The object, a $6\mathrm{cm}$ cube, is unconstrained. The hand’s wrist is welded to the world frame. The task is to reorient the cube to target $SE(3)$ poses.
V-A Experiment Setup
V-A1 Goal Generation
@@
For both the IiwaBimanual and AllegroHand systems, we generate about 1000 pairs of initial conditions and goals which are locally reachable yet far enough to be challenging for MPC. Details about the goal generation scheme can be found in Section X-C.
@@
V-A2 Evaluation Metrics
We evaluate the performance of MPC by comparing the difference between $q^{\mathrm{o}}_{\text{final}}$, the final object configuration reached by the by Algorithm 2, and $q^{\mathrm{o}}_{\text{goal}}$, the goal object configuration. As we are only concerned with reaching object goals, we set the robot-related entries in the cost matrix $\mathbf{Q}$ in (37) to 00, and rely on the action cost (37b) and the CTR constraint (37d) to stay close to the linerization points.
We split the object configuration ${q^{\mathrm{o}}}$ into a quaternion $Q$ and a position $p$: ${q^{\mathrm{o}}}\coloneqq(Q,p)$, both expressed relative to the world frame. We divide the error in ${q^{\mathrm{o}}}$ into a translation error $\|p_{\text{goal}}-p_{\text{final}}\|$, and an rotation error $\Delta\theta(Q_{\text{goal}},Q_{\text{final}})$, where $\Delta\theta(\cdot,\cdot)$ returns the angular difference between two unit quaternions.
V-A3 Implementation Details
The numerical experiments are run on a M2 Max Macbook Pro with 64GB of RAM. We use the same open-sourced implementation of the CQDC dyanmics as in [11]. The convex subproblem (37) in Algorithm 1 is solved with Mosek [66].
@@ V-A3 Implementation Details
Figure 9: Results for running Algorithm 2 with R-CTR, CTR and ETR constraints on the pairs of goals and initial conditions generated in Section V-A1. The histogram shows how many samples reached a specified error in the $x$ axis, where error is defined as $\|p_{\text{goal}}-p_{\text{final}}\|+|\Delta\theta(Q_{\text{goal}},Q_{\text{final}})|$. The dotted lines in the histogram represent the average errors. Some dotted lines are barely visible due to overlapping (their values are available in Table II). Each dot in the scatter plot corresponds to a pair of an initial condition $q^{\mathrm{o}}_{0}$ and a goal $q^{\mathrm{o}}_{\text{goal}}$. Color of the dots is determined by the error: warmer colors indicate bigger errors. The three scatter plots in each row share the same color scale, whose maximum value is the average of the errors for reaching the goals with ETR. (a) For IiwaBimanual, the $x$ axis of the scatter plots is the position difference between $q^{\mathrm{o}}_{0}$ and $q^{\mathrm{o}}_{\text{goal}}$, the $y$ axis the angular difference. (b) For AllegroHand, as the position difference is small for all goals, the $x$ axis shows the yaw angle difference between $q^{\mathrm{o}}_{0}$ and $q^{\mathrm{o}}_{\text{goal}}$, and the $y$ axis the pitch angle difference. The roll, pitch and yaw axes are defined in Figure 14a.
V-B Effect of Trust Region Radius $\mathbf{\Sigma}$ and Rollout Horizon $H$
In this section, we explore how the two parameters affect the performance of our MPC scheme on both the IiwaBimanual and AllegroHand systems. We also study how these parameters interplay with different trust region formulations, including R-CTR, CTR and ETR. We measure performance in terms of average tracking errors for the goals selected in Section V-A1.
As we sweep the two parameters, we fix the planning horizon $T=1$ in Algorithm 2, since increasing $T$ does not improve performance (see Section X-D for details). Furthermore, we use $n_{\mathrm{max}}=2$ IiwaBimanual and $n_{\mathrm{max}}=3$ for AllegroHand.
V-B1 Trust Region Radius
How does the trust region radius $\mathbf{\Sigma}$ affect the performance of R-CTR, CTR and ETR? To answer this question, we run Algorithm 2 on both systems with different trust region radii. We set $\mathbf{\Sigma}$ to a constant scaling of the identity matrix, i.e. $\mathbf{\Sigma}=r^{-2}\mathbf{I}$ where $r\in\mathbb{R}_{+}$ is the radius. The results are illustrated in Figure 10a and b, where we measure performance by the average tracking errors for the goals generated in Section V-A1.
Figure 10: How tracking errors, defined as the sum of translation and rotation errors, change with different trust region radius $r$ ($\mathbf{\Sigma}=r^{-2}\mathbf{I}$) and rollout horizon $H$. The plotted curves denote the average error for the goals generated in Section V-A1, and the error bars denote $0.1\sigma$ where $\sigma$ is the standard deviation.
V-B2 Rollout Horizon $H$
V-C Can Our Method Reach Goals?
Finally, for all initial condition and goal pairs generated in Section V-A1, we run Algorithm 2 and analyze the results using the metrics in Section V-A2. A subset of the MPC hyperparameters is summarized in Table I.
$T$ $n_{\mathrm{max}}$ $r$ $H$
IiwaBimanual 1 2 0.1 20
@@ V-C Can Our Method Reach Goals?
TABLE I: MPC hyperparameters: $T$ and $n_{\mathrm{max}}$ are defined in Algorithm 1; $r$ is the trust region radius ($\mathbf{\Sigma}=r^{-2}\mathbf{I}$); $H$ is the MPC rollout horizon in Algorithm 2. These parameters are chosen based on Figure 10.
As shown in Figure 9, MPC with R-CTR achieves the lowest average error on both IiwaBimanual and AllegroHand. As Figure 9 combines rotation and translation errors, we present them separately in Table II to better illustrate their relative magnitudes. Notably, MPC with R-CTR also achieves the lowest variance.
@@
Trans. [mm] Rot. [mrad] Tran. [mm] Rot. [mrad]
@@
TABLE II: Mean translation and rotation errors for the combined errors in Figure 9. Each cell displays the mean (standard deviation).
On IiwaBimanual, R-CTR does significantly better than CTR and ETR. In contrast, for AllegroHand, the advantage of R-CTR over the other variants is much more nuanced. This corroborates our analysis in Section V-B1.
We also report the wall clock time for the most time-consuming steps in Algorithm 2, shown in Table III. For AllegroHand, $\mathbf{CtrTrajOpt}$ is considerably slower due to its high-dimensional state and action space, as well as the larger number of collision pairs. The slowness is further compounded by our CQDC dynamics implementation, which has not yet been optimized to leverage problem sparsity. Among the three trust region variants, ETR is the fastest due to the absence of feasibility constraints. In contrast, CTR is the slowest, as it enforces both primal and dual feasibility.

### v3

@@ 8 $n\leftarrow n+1$
Algorithm 1 CTR Trajectory Optimization-CtrTrajOpt
4.2 Initial Guess Heuristic
As the trajectory optimization problem (36) is nonconvex, the solution that Algorithm 1 converges to can be sensitive to the choice of initial guesses for the input trajectory $\bar{u}_{0:T-1}$. The most uninformed initial guess would be to set the position command to be the current configuration of actuated bodies ($\bar{u}_{t}=\bar{q}^{\mathrm{a}}_{0},\forall t$). However, this can be problematic especially when the robots are not in contact with the object in $q_{0}$, as $\mathbf{B}_{\kappa}^{\mathrm{o}}$, the gradient of object motion w.r.t. robot actions, can get close to 0 even with dynamics smoothing when the distance between them are far away.
@@ 4.2 Initial Guess Heuristic
and simulate forward using Drake (Han et al., 2023) until the robot is in contact with the object. We use a generous amount of smoothing (e.g. $\kappa\in[10,100]$) to ensure that the contact forces $\lambda_{i}$ are large enough to pull the robot links towards the object even if they are far away at $q_{0}$.
4.3 Examples
@@
Applying the initial guess heuristics in Section 4.2 brings the robot into contact with the object, yielding $^{1}\bar{u}=0.0$ (magenta square in Figure 6). The updated $^{1}\bar{u}$ produces the new linearized dynamics (magenta line in Figure 6a), whose slope more closely matches the “in contact” piece of the non-smooth contact dynamics. Meanwhile, the dual feasibility constraint (27d) keeps the ball from pulling the box back.
@@
  • Red zone: If $q^{\mathrm{o}}_{\text{goal}}$ is in the red zone, the optimizer would move backwards but stop at the boundary of the red region, which corresponds to the leftmost point in the magenta line segments in Figure 6. Without dual feasibility (27d), a naive linear model with an ETR will move backwards into the red region to attempt pulling the object.
@@
Lastly, we plot the A-CTR and the RA-CTR for the above sequence of $\bar{u}$ in Figure 7e. Here we use $\mathbf{\Sigma}=\mathrm{diag}(0.05^{-2},0.05^{-2})$, which bounds $\delta u$ in a circle of radius $0.05$. The curved part of the trust region boundaries arises from the ellipsoidal constraint (25c), whereas the straight part follows from the feasibility constraints (25d) and (25e). Compared with the RA-CTR at $^{0}\bar{u}$ (dark yellow), the RA-CTR at $^{1}\bar{u}$ (magenta), obtained from applying the initial guess heuristics to $^{0}\bar{u}$, approximates the sticking region much better. Furthermore, as $^{2}\bar{u}$ (dark green) gets deeper into the “sticking” mode, the straight edge of its RA-CTR aligns more closely with the boundary between sticking and sliding. Finally, compared to the RA-CTRs, the non-relaxed A-CTRs are always smaller. As a result, Algorithm 1 would need more iterations to converge if the RA-CTR constraint in (37d) were replaced by A-CTRs.
4.4 Model Predictive Control
The CTR-based trajectory optimizer in Section 4.1 can be readily turned to a controller through Model Predictive Control (MPC). As shown in Algorithm 2, we obtain the optimal trajectory $u^{\star}_{t}$ by solving (36) with Algorithm 1 (Algorithm 2), only use the first input $u^{\star}_{0}$ to deploy on the true CQDC dynamics Algorithm 2, then re-plan after observing the next configuration. Moreover, we initialize $\mathbf{CtrTrajOpt}$ from the initial guess heuristics (Section 4.2) at the first iteration (Algorithm 2), and from the solution of the previous iteration at later iterations (Algorithm 2). We refer to the length of the MPC episode as $H$, as opposed to planning horizon $T$.
@@ 13 $q_{t+1}=f(q_{t},u_{0}^{\star})$;
5 Local MPC Experiments
In this section, we evaluate the performance of the CTR-based MPC in Algorithm 2 under the CQDC dynamics constraints. We are particularly interested in answering the following questions:
@@ 5 Local MPC Experiments
  • the planar IiwaBimanual system (Figure 9-top and Example 6), which comprises 3 unactauted DOFs, 6 actuated DOFs and 29 collision geometries. Both the arms and the object are constrained to move in the $xy$ plane, with gravity pointing along the negative $z$ direction. The object is a cylinder measuring $0.28\mathrm{m}$ in diameter. The task is to rotate the object to target $SE(2)$ poses. Each robot arm’s collision geometries consist of 14 spheres; the object collision geometry is a single cylinder.
  • the 3D AllegroHand system (Figure 9-bottom), which comprises 6 unactauted DOFs, 16 actuated DOFs and 39 collision geometries. The object, a $6\mathrm{cm}$ cube, is unconstrained. The hand’s wrist is welded to the world frame. The task is to reorient the cube to target $SE(3)$ poses. The hand has 30 collision geometries, which consist of a collection of boxes, spheres and cylinders; the object collision geometries include a cube and 8 small spheres at the cube’s corners.
Figure 9: Visual and collision geometries of IiwaBimanual (top) AllegroHand (bottom). Every IiwaBimanual geometry is shown as white, while the robot collision geometries in AllegroHand are colored to increase contrast.
5.1 Experiment Setup
5.1.1 Goal Generation
@@
For both the IiwaBimanual and AllegroHand systems, we generate about 1000 pairs of initial conditions and goals which are locally reachable yet far enough to be challenging for MPC. Details about the goal generation scheme can be found in Section A.3.
@@
5.1.2 Evaluation Metrics
We evaluate the performance of MPC by comparing the difference between $q^{\mathrm{o}}_{\text{final}}$, the final object configuration reached by the by Algorithm 2, and $q^{\mathrm{o}}_{\text{goal}}$, the goal object configuration. We split the object configuration $q^{\mathrm{o}}$} into a quaternion $Q$ and a position $p$: ${q^{\mathrm{o}}}\coloneqq(Q,p)$, both expressed relative to the world frame. We divide the error in $q^{\mathrm{o}}$} into a translation error $\|p_{\text{goal}}-p_{\text{final}}\|$, and an rotation error $\Delta\theta(Q_{\text{goal}},Q_{\text{final}})$, where $\Delta\theta(\cdot,\cdot)$ returns the angular difference between two unit quaternions.
5.1.3 Implementation Details
The numerical experiments are run on a M2 Max Macbook Pro with 64GB of RAM. We use the same open-sourced implementation of the CQDC dynamics as in Pang et al. (2023). The convex subproblem (37) in Algorithm 1 is solved with Mosek ApS (2024).
5.2 Effect of Trust Region Radius $\mathbf{\Sigma}$ and Rollout Horizon $H$
In this section, we explore how the two parameters affect the performance of our MPC scheme on both the IiwaBimanual and AllegroHand systems. We also study how these parameters interplay with different trust region formulations, including R-CTR, CTR and ETR. We measure performance in terms of average tracking errors for the goals selected in Section 5.1.1.
As we sweep the two parameters, we fix the planning horizon $T=1$ in Algorithm 2, since increasing $T$ does not improve performance (see Section A.4 for details). Furthermore, we use $n_{\mathrm{max}}=2$ IiwaBimanual and $n_{\mathrm{max}}=3$ for AllegroHand.
5.2.1 Trust Region Radius
How does the trust region radius $\mathbf{\Sigma}$ affect the performance of R-CTR, CTR and ETR? To answer this question, we run Algorithm 2 on both systems with different trust region radii. We set $\mathbf{\Sigma}$ to a constant scaling of the identity matrix, i.e. $\mathbf{\Sigma}=r^{-2}\mathbf{I}$ where $r\in\mathbb{R}_{+}$ is the radius. The results are illustrated in Figure 10a and b, where we measure performance by the average tracking errors for the goals generated in Section 5.1.1.
Figure 10: How tracking errors, defined as the sum of translation and rotation errors, change with different trust region radius $r$ ($\mathbf{\Sigma}=r^{-2}\mathbf{I}$) and rollout horizon $H$. The plotted curves denote the average error for the goals generated in Section 5.1.1, and the error bars denote $0.1\sigma$ where $\sigma$ is the standard deviation. For plots (e) and (f), the plotted curves denote the best average error across all combinations of $r$ and $H$ with constant $r\cdot H$.
5.2.2 Rollout Horizon $H$
5.2.3 $r\cdot H$: “Effective Lookahead”
The product of the trust region radius $r$ and the planning horizon $H$ represents the effective lookahead of MPC. A short, aggressive plan (low $H$, high $r$) might change the system state as much as a long, cautious one (high $H$, low $r$), giving them a similar reach into the future state of the system. For a fixed $r\cdot H$, we expect a more “informative” trust region can better guide the system towards the goals, resulting in lower final tracking errors.
In Figure 10(e) and (f), we examine how the errors change as a function of the effective lookahead. To fairly measure the average error at a fixed value of $r\cdot H$, each data point on these plots represents the best performance achieved for a given $r\cdot H$ value, found by varying $r$ and $H$ while holding $r\cdot H$ constant.
For sufficiently large $r\cdot H$, R-CTR and CTR significantly outperform ETR, suggesting that the incorporation of the feasibility constraints provides better guidance to MPC, even when ETR is used cautiously (low $r$ and high $H$).
@@ 5.2.3 $r\cdot H$: “Effective Lookahead”
Figure 11: Results for running Algorithm 2 with R-CTR, CTR and ETR constraints on the pairs of goals and initial conditions generated in Section 5.1.1. The histogram shows how many samples reached a specified error in the $x$ axis, where error is defined as $\|p_{\text{goal}}-p_{\text{final}}\|+|\Delta\theta(Q_{\text{goal}},Q_{\text{final}})|$. The dotted lines in the histogram represent the average errors. Some dotted lines are barely visible due to overlapping (their values are available in Table 3). Each dot in the scatter plot corresponds to a pair of an initial condition $q^{\mathrm{o}}_{0}$ and a goal $q^{\mathrm{o}}_{\text{goal}}$. Color of the dots is determined by the error: warmer colors indicate bigger errors. The three scatter plots in each row share the same color scale, whose maximum value is the average of the errors for reaching the goals with ETR. (a) For IiwaBimanual, the $x$ axis of the scatter plots is the position difference between $q^{\mathrm{o}}_{0}$ and $q^{\mathrm{o}}_{\text{goal}}$, the $y$ axis the angular difference. (b) For AllegroHand, as the position difference is small for all goals, the $x$ axis shows the yaw angle difference between $q^{\mathrm{o}}_{0}$ and $q^{\mathrm{o}}_{\text{goal}}$, and the $y$ axis the pitch angle difference. The roll, pitch and yaw axes are defined in Figure 15a.
5.3 Can Our Method Reach Goals?
Finally, for all initial condition and goal pairs generated in Section 5.1.1, we run Algorithm 2 and analyze the results using the metrics in Section 5.1.2.
Table 2 summarizes a subset of MPC hyperparameters which are chosen based on the discussion in Section 5.2. Additional hyperparameters are summarized in Section A.5.
System $T$ $n_{\max}$ $r$ $H$
IiwaBimanual 1 2 0.10 20
@@ 5.3 Can Our Method Reach Goals?
Table 2: MPC hyperparameters: $T$ and $n_{\mathrm{max}}$ are defined in Algorithm 1; $r$ is the trust region radius ($\mathbf{\Sigma}=r^{-2}\mathbf{I}$); $H$ is the MPC rollout horizon in Algorithm 2.
5.3.1 Goal Tracking Performance
As shown in Figure 11, MPC with R-CTR achieves the lowest average error on both IiwaBimanual and AllegroHand. As Figure 11 combines rotation and translation errors, we present them separately in Table 3 to better illustrate their relative magnitudes. Notably, MPC with R-CTR also achieves the lowest variance.
@@ 5.3.1 Goal Tracking Performance
Trans. [mm] Rot. [mrad] Trans. [mm] Rot. [mrad]
@@
Table 3: Mean translation and rotation errors for the combined errors in Figure 11. Each cell displays the mean (std).
Upon a closer examination of the distribution of errors in Figure 11, a key advantage of R-CTR is that it produces far fewer trajectories with larger errors. For IiwaBimanual, this advantage is significant: the R-CTR distribution is more tightly clustered at low error values and lacks the heavy tail in CTR and ETR’s error distributions. For AllegroHand, the result is more nuanced: R-CTR and CTR perform almost identically, but both are a small yet discernible shift towards the left compared to ETR.
We believe the benefit of R-CTR is the most significant when the dual feasibility constraints substantially restrict their corresponding larger ETR. As indicated by the examples in Section 3, this restriction is pronounced in the unilateral contact scenarios (Example 1 and Example 3), but subtle or even non-existent when the contact points form a stable, force-closure grasp (e.g. Example 2).
This hypothesis is corroborated by the example trajectories in Section A.8. In the ETR failure cases (Figure 24a and Figure 25a), the robots lack a force-closure grasp and cannot move the object in certain directions. The dual feasibility constraint in R-CTR correctly identifies these limitations and prevents the extraneous actions that cause the ETR baseline to fail. In contrast, when the robots have an effective force-closure grasp on the object (Figure 24b and Figure 25b), including the dual feasibility constraints is less crucial for success.
Another key question is why the primal feasibility constraint makes CTR overly restrictive on IiwaBimanual (hence the larger tracking errors), while having almost no negative effect on AllegroHand when compared to R-CTR. We hypothesize that the impact of this constraint is highly dependent on the system’s kinematics and the specific contact configuration. As our examples illustrate, the primal feasibility constraint can sometimes severely reduce the trust region volume (e.g. Figure 7e) but at other times barely alters it (e.g. the blue nominal configuration in Figure 4). This suggests the grasping configurations on AllegroHand are less affected by this particular constraint. A deeper analysis of the role of primal feasibility in different contact scenarios could be an interesting topic for future research.
5.3.2 Infeasibility
Encountering infeasible constraints is a common issue that plagues iterative methods for solving the non-convex trajectory optimization problem (36). However, Algorithm 2 rarely suffers from this problem. As shown in Table 4, despite each goal requiring the computation of tens ($T\cdot H\cdot n_{\mathrm{max}}$) of unique trust regions, the empirical rate of infeasibility is very low. As a result, when generating the tracking results in Figure 11, we simply terminate MPC prematurely if an infeasible trust region is encountered.
ETR CTR R-CTR Number of Goals
IiwaBimanual 0 0 0 1233
AllegroHand 5 2 3 1000
Table 4: Number of infeasible trust regions encountered while running Algorithm 2 to reach the goals from Section 5.1.1.
5.3.3 Computation Time
We report wall clock time for the most time-consuming steps in Algorithm 2, shown in Table 5. For AllegroHand, $\mathbf{CtrTrajOpt}$ is slower due to its high-dimensional state and action space, as well as larger number of collision pairs. Among the three trust region variants, ETR is the fastest due to the absence of feasibility constraints. In contrast, CTR is the slowest, as it enforces both primal and dual feasibility.
There is significant room for improvement for the computation time, which we leave as future work. For instance, we can (i) exploit common sparsity patterns in MPC programs (Mastalli et al., 2020), and (ii) prune the trust region constraints that are too far from making contact (our current trust region construction is very conservative, see Section A.5 for a more detailed discussion).